OpenStudy (anonymous):

What is the solution to the equation 9 −3x ≈ 7?

undeadknight26 (undeadknight26):

hmm is that a fancy = sign or something else?

undeadknight26 (undeadknight26):

if the equation is: 9 - 3x = 7 +9 +9 3x = 16 /3 /3 x = 5.33

OpenStudy (anonymous):

x = 0.376 x = 0.295 x = −0.295 x = −0.376

OpenStudy (hitaro9):

-3x = 7-9 -3x = -2 x = -2/-3 x= 2/3

undeadknight26 (undeadknight26):

that would equal .6666 right...hmmm

OpenStudy (hitaro9):

Are those the only options woohoo?

OpenStudy (anonymous):

yeah these are the only options

OpenStudy (hitaro9):

Could you have made a typo in the original problem? Cause I'm not seeing where I could have made a mistake?

OpenStudy (anonymous):

What is the solution to the equation 9 ^−3x ≈ 7?

OpenStudy (hitaro9):

Oh

OpenStudy (hitaro9):

That's a completely different problem lol

OpenStudy (anonymous):

@undeadknight26 that weird equal sign is used when you are making an approximation.

undeadknight26 (undeadknight26):

ohhhh....

OpenStudy (hitaro9):

Its gunna be -.295

OpenStudy (anonymous):

Thank you but can you explain how you got it?

undeadknight26 (undeadknight26):

yes yes it is...i didn't see the ^ sign...

OpenStudy (hitaro9):

Take the log of both sides -3x log 9 = log 7

undeadknight26 (undeadknight26):

this is stuff i do not know so umm im just gonna you know..

OpenStudy (hitaro9):

x = log7/-3log9

OpenStudy (hitaro9):

And when you plug that in the calculator, its -.295

OpenStudy (hitaro9):

I dunno how'd you figure it out without a calculator, but I'm assuming you're allowed to use one